OpenStudy (anonymous):

a body travelling in straight line travels 2m in the first 2 seconds and 2.2m in the next four second with constant retardation. What will be its velocity at the end of seventh second from the start

OpenStudy (anonymous):

plzzanswer

OpenStudy (anonymous):

Let me solve it myself please)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Formula for motion with constant acceleration is in one dimension, which is the case s = s0+v0t+at^2/2 You can definitely find what you need using it

OpenStudy (anonymous):

Try to solve it yourself

OpenStudy (anonymous):

elaborate

OpenStudy (anonymous):

Ask me what you don't understand in how to use it.

OpenStudy (anonymous):

i understood nothin

OpenStudy (anonymous):

Ok. It's better for you in future to read about kinematics. So what is given is t1 = 2 s, s1 = 2 m t2 = 6 s, s2 = 4.2 m s0 = 0 because we arranged our system of coordinates so that initial position of an object at the origin of it. a - which is acceleration, is constant, so we can use equations of kinematics for motion with constant acceleration. Do you get that?

OpenStudy (anonymous):

Is that clear, what I wrote?

OpenStudy (anonymous):

Hey man are you here?;)

OpenStudy (anonymous):

i got it but i don't know kinematics

OpenStudy (anonymous):

Ok, it doesn't matter. Now our equation becomes s = v0t+at^2/2 using values of s1 and t1 you can express v0 in term of a.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You can type in your current data so we can check if it is right.

OpenStudy (anonymous):

v0 - ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No) try again

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Ok I'll show you the steps

OpenStudy (anonymous):

s1 = 2m, t1 =2s 2 = v0 *2 + a 2^2/2 1 = v0 + a v0 = 1 - a

OpenStudy (anonymous):

and our equation becomes more "simple" s = (1-a)t + at^2/2

OpenStudy (anonymous):

it should be 1-2a i think?

OpenStudy (anonymous):

Ok let's try again 2 = v0 *2 + a 2^2/2 2 = v0 *2 + a *4/2 2 = v0 *2 + a *2 1 = v0 + a v0 = 1 - a Do you have any doubts about this now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am going on: and our equation becomes more "simple" s = (1-a)t + at^2/2

OpenStudy (anonymous):

yeah iam gettin it now

OpenStudy (anonymous):

Putting in s and t values you can find a. When you will find a, using another equation of kinematics for motion with constant acceleration, you can easily find v for any moment of time v = v0 + at

OpenStudy (anonymous):

Notice that you will get negative value for acceleration because its retardation, if it's not negative, be sure you got wrong value.

OpenStudy (anonymous):

i know these equation v=u+at s=ut+1/2at^2 v^2-U^2=2as are these equation of kinematics

OpenStudy (anonymous):

I'm not sure about the last one because it takes time to derive it, yes these are equations of kinematics for motion with constant acceleration.

OpenStudy (anonymous):

Here U is initial velocity!

OpenStudy (anonymous):

i know